The current in a long solenoid of 95 turns generates a magnetic flux within it of 65.0 mWb per turn. If that current is gradually reversed in a time of 170. ms, what will be the induced emf?
Answer #1

If we assume that all the positive directions have been assigned in a consistent way, then the induced emf will be positive, as the change in flux is negative (+65.0 mWb/turn to -65.0 mWb/turn, a change of -130 mWb/turn).


emf = -N*d(change in flux) /dt

emf = -95*(-130x10^-3/170x10^-3 =72.647 V